OpenStudy (nathanjhw):

The change in the momentum of an object (Δ p) is given by the force, F, acting on the object multiplied by the time interval that the force was acting: Δ p = F Δt . If the force (in newtons) acting on a particular object is given by F(t)=cost , what’s the total change in momentum of the object from time t = 5 until t = 7 seconds?

OpenStudy (nathanjhw):

0.402 newton•sec 0.708 newton•sec 0.909 newton•sec 1.416 newton•sec 1.616 newton•sec

OpenStudy (anonymous):

0.909 newton sec

OpenStudy (nathanjhw):

How did you solve it?

OpenStudy (nathanjhw):

@TheSmartOne

TheSmartOne (thesmartone):

I don't know calculus... ._. But maybe this can help you :) https://answers.yahoo.com/question/index?qid=20120426232252AAW7OZN http://mathhelpforum.com/calculus/226439-momentum-question.html

OpenStudy (perl):

you can use the integral

OpenStudy (nathanjhw):

The integral of?

OpenStudy (perl):

$$ \Large{ \Delta p = F\Delta t\\ \therefore\\ Total~ Momentum = \sum \Delta p = \sum F \Delta t \\ \therefore \\ \int dp = \int F(t) dt = \int_{5}^{7} \cos t dt } $$

OpenStudy (nathanjhw):

\[\Delta p = \sin(7) - \sin(5)\]

OpenStudy (nathanjhw):

correct?

OpenStudy (perl):

correct

OpenStudy (nathanjhw):

That equals 0.0347, and that's not one of my answers.

OpenStudy (perl):

you are using degree mode

OpenStudy (perl):

not radian mode, the question is in radians

OpenStudy (nathanjhw):

Well that explains a lot. I got 1.616.

OpenStudy (nathanjhw):

Thank you!

OpenStudy (nathanjhw):

@perl

OpenStudy (perl):

your welcome